Rick’s Department Stores has had the following pattern of earnings per share over the last five years:  

Year Earnings
per share
20XU $ 14.00  
20XV   14.70  
20XW   15.44  
20XX   16.21  
20XY   17.02  
 

The earnings per share have grown at a constant rate (on a rounded basis) and will continue to do so in the future. Dividends represent 40 percent of earnings.

  

a. Project earnings and dividends for the next year (20XZ). (Do not round intermediate calculations. Round the final answers to 2 decimal places.)

  

  20XZ
  Earnings $  
Dividend $  
 

  

b. If the required rate of return is 13 percent, what is the anticipated share price at the beginning of 20XZ? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  

Anticipated stock price           $
Answer #1
Req 1.
Earnings growith rate = (14.70-14) / 14 *100 = 5%
Project Earnings for next year
Earnings for last year 17.02
Add: Growth @5% 0.85
Earnings for next year 17.87
Dividend for next year:
Earnings for nxt year 17.87
Multiply: Dividend payout % 40%
Dividend for next year: 7.15
Req 2.
Stock price = Dividend of next year / (Required rate -Growth rate)
7.15 / (13-5)% = 89.38
Stock price = $89.38